Subject: 3.5-m Users Committee minutes, 1/12/09

From: Suzanne Hawley

Submitted: Tue, 20 Jan 2009 23:44:59 -0800 (PST)

Message number: 1124 (previous: 1123, next: 1125 up: Index)

                APO 3.5-m Users Committee Phonecon, 1/12/09

Attending:  Suzanne Hawley, Scott Anderson, Mark Klaene, Bill 
Ketzeback, Russet McMillan, John Bally, Remy Indebetouw, Bruce 
Gillespie, and Michael Strauss

Absent: Al Harper, Jon Holtzman, and Jon Fulbright


User feedback, comments from institutional representatives:

Princeton (Michael Strauss) - Michael said things were fine, with no 

Colorado (John Bally) - John reported that users had recent problems 
with pointing and slit-angle problems using TripleSpec.  Mark replied 
that the site staff were working on the problems with telescope 
pointing and the rotator.  John advised that TripleSpec users should 
come prepared with wide-field pointing charts until the pointing and 
rotator problems are resolved.

Johns Hopkins (Jon Fulbright) - no report.

Washington (Scott Anderson) - Scott has nothing to report.  Suzanne 
added a general thank-you to Min-Su Shin (Princeton) for graciously 
accommodating a telescope time change last month.

Chicago (Al Harper) - no report.

New Mexico State (Jon Holtzman) - no report.

Virginia (Remy Indebetouw) - Remy said his users were mostly happy. 
One user asked why we delete old data, given the low cost 
of disk storage these days.  Suzanne said that it has long been our 
policy that APO is not a data archive because we cannot afford the
time and effort.  She reminded everyone that the current policy is
that data are maintained on newton for between 9-12 months (they
are deleted at the beginning of the same quarter of the following
year, e.g. Q1 2008 data were deleted on Jan 1, 2009). 
Because of the large number of images obtained with the 
recommended Fowler Sampling on NIC-FPS, the raw NIC-FPS frames 
are kept only for three months.  The processed NIC-FPS images
are kept for the normal 9-12 months.


Discussion of telescope/instruments report - Klaene:

Mark summarized the report, spotlighting the recent issues with the 
tertiary axial actuators.  Also, a problem with the instrument rotator 
was worked on and appears to be resolved.  The DIS-blue ion 
pump was replaced for a second time; this time it seems to be working 
well.  Also, the DIS slit-viewer shutter was replaced twice, and is 
now working reliably.  The other instruments are performing nominally, 
and although mounting TripleSpec is somewhat improved, it is still 
more difficult than the other instruments.  The new rotator for Agile 
is mounted at the TR2 port and being tested. We plan to move the 
instrument in February.  The CDR for the telescope axes direct-drive 
system will be in early March, and installation is planned for 
the summer shutdown.


3.5-m Telescope, Instruments, and CIF Projects Highlights, 12/2/08 
through 1/07/09

                                 Mark Klaene

0) Overview

Winter has started to set in with a number of snowfalls and mostly 
subfreezing temperatures.  Telescope problems were generally 
associated with the tertiary, however these were much reduced from 
previous months.  The holiday came and went with minimal disruptions 
to the staff's holiday leave schedule.  Site infrastructure held its 
own throughout the holiday break. The observatory now has a video 
conferencing system on site.

1) Telescope

NA2 axis controller has been performing very well.

Tertiary rotation: rotation is working well, we had one instance of 
the axial actuator encoders failing. The problem went away on its own, 
which is not something we wanted to see happen.

We are seeing some pointing errors that seem to be related to tertiary 
rotation but our telemetry (which is quite extensive) does not 
indicate a problem.  This is still being pursued.

2) Instruments

The DIS-blue ion pump was replaced again and is now working well.  The 
DIS slitviewer shutter went through two replacements and we are 
waiting for further testing to verify the second one is working.

DIS and SPIcam are both fully operational.

Echelle and NIC-FPS both continue to be operational with no problems. 
We are still waiting for UVa to design and build up a new NIC-FPS fan- 
out board that will hopefully address the residual readout noise.

TripleSpec mounting is still occasionally giving us some problems, but 
recent modifications and more training seem to have improved the 

Agile is still available at NA2; it will be moved to the TR2 port once 
the new rotator is fully tested.  The rotator has been installed and 
is ready for testing.  TUI integration for Agile continues.  Some
software integration issues are being worked.

APOLLO remains operational.  The high-precision gravimeter is now 
slated for arrival in January, and preparations are in progress for 
its installation at the 3.5-m.

3) CIF projects

A date of March 2nd as been set for the direct-drive CDR.


Update on Agile rotator and TUI integration - Klaene, Hawley

The TUI integration of Agile is proceeding well, and some on-sky 
testing has been accomplished.  There are uncertainties about the 
accuracy of the timing, which will soon be tested, but in the meantime 
the TUI interface is working well enough for user test-drives if high 
accuracy timing is not required.

As previously discussed at an earlier meeting, there is no guider 
planned for the new mount position for Agile because of space 
limitations and there being no compelling science reason to guide its 
short exposures.  However, it has been suggested that guiding may be 
desirable to get the highest photometric precision by keeping the 
field registered to the same pixels.  It is possible to use the 
science data itself to guide the telescope if necessary for precision 
flat-fielding, but this capability is not currently planned for 
implementation.  Suzanne asked that committee members make sure that 
their Agile users know that there will be no guiding of Agile 
exposures starting in mid-February, and if there are users that feel
guiding is needed, they should bring the issue to Suzanne and Mark.


Update on Triplespec scheduling during Q1 2009 - Hawley

Suzanne reported that TripleSpec is partially block-scheduled in Q1, 
and off-block usage in December and Q1 will help us determine how 
feasible it is to allow its general scheduling.  The mounting 
procedure for the instrument is still difficult, but improvements have 
been made and more are in the works.  Users can ask to use TripleSpec 
in off-block-scheduled times, but they need to be aware that the 
instrument mounting is more time-consuming than for the other
instruments, and to be prepared to be patient.


Instrument studies - Hawley

We have received expressions of interest for several new or upgraded 
instruments.  Cynthia Froning (Colorado) is looking at wide-field 
visible imaging capabilities, Eric Burgh (Colorado) is looking 
at upgrades for the echelle, and Jian Ge (Florida) is proposing an
NIR dispersed interferometer for RV studies and a high resolution
NIR spectrograph.  If any other users are interested in these or other 
instrumentation projects, they should contact the aforementioned or 
Suzanne and Mark.


New items

Bill Ketzeback said that we need to reevaluate the existing user 
information for NIC-FPS, some of which may be different since it was 
first posted because the read electronics for the detector have been 
changed from 2-channel to 16-channel reads.  This may have affected 
the linearity and read noise.  John Bally said that these values
should be updated, but he will wait until the new fan-out board is 
installed.  Mark added that the NIC-FPS sub-frame readout times and 
other related details also need to be updated.



[open from previous months]:


[new actions from last meeting]:

ACTION:  Focus offsets for the various NIC-FPS filter need to be 
determined, documented for the users, and automated corrections 
implemented in the operations software.

     STATUS:  John Bally said they have received engineering data 
from APO, and the filters are definitely not parfocal and 
there are significant offsets in their focal distances. 
Data in better seeing, or an alternative method (e.g., 
Hartmann or phase-retrieval tests) would be helpful in order to better 
quantify the focus shifts between filters.  The measured focus shifts 
need to be added to the user information for NIC-FPS, and users are 
warned to focus each filter separately.  The focus shifts are in the 
25 to 75 micron range.  It is possible that the ICC software could 
automatically make the focus correction, but this would require
significant effort to implement.

ACTION:  Anybody with an interest in leading and/or supporting new/ 
upgraded instruments for the telescope should make their intentions 
known to Suzanne Hawley by next month's Users Committee meeting.

     STATUS:  Closed, see discussion this meeting.


Next meeting: The next Users Committee phonecon will be on Monday, 2 
February, at 8:30 AM Pacific Time.  The agenda and other materials 
will be sent to the committee members during the preceding week.

APO APO APO APO APO  Apache Point Observatory 3.5m  APO APO APO
APO  This is message 1124 in the apo35-general archive. You can find
APO  the archive on
APO  To join/leave the list, send mail to
APO  To post a message, mail it to